OpenStudy (anonymous):

Never solved this type of thing before, am I right? Solve the following equation. If necessary, enter your answer as an expression involving natural logarithms or as a decimal approximation that is correct to at least four decimal places. 2^(2x+6)=3^(x-46) x = 0.9086 (don't think I am).

OpenStudy (mathstudent55):

Take the log of both sides.

OpenStudy (anonymous):

\[ 2^{2x+6}=3^{x-46} \] \[(2x+6)\ln(2)=(x+46)\ln(3)\] and a raft of algebra to solve for \(x\)